OpenStudy (aayushi.somani):

integrate by substitution: (e^x + 1/e^x) ^2 please show all the steps..

OpenStudy (solomonzelman):

You really have to do it by substitution? because I had an easier set of steps. 1) re-write the 1/e^x as e^(-x). 2) use (a+b)²=a²+2ab+b² 3) Integrate the sum term by term optional 4) After you get every you can factor out of 1/2.

OpenStudy (solomonzelman):

after you get every *integral

OpenStudy (aayushi.somani):

actually its the ques to do it by substitution.. please help me through substitution.. @SolomonZelman

OpenStudy (solomonzelman):

\[\int\limits_{ }^{ } \left(\begin{matrix} e^x+\frac{1}{e^x} \end{matrix}\right)^2~dx\]\[\int\limits_{ }^{ } \left(\begin{matrix} e^x+e^{-x} \end{matrix}\right)^2 ~dx\]\[\int\limits_{ }^{ } \left(\begin{matrix} e^{-x}+e^x \end{matrix}\right)^2 ~dx\]\[\int\limits_{ }^{ } \left(\begin{matrix} e^{-2x}+e^{2x}+2 \end{matrix}\right) ~dx\]\[\int\limits_{ }^{ } e^{-2x}~dx+\int\limits_{ }^{ } e^{2x}~dx+\int\limits_{ }^{ } 2~dx \]

OpenStudy (solomonzelman):

I am not very sure what and where to sub in this case... and esp, why

OpenStudy (aayushi.somani):

ohk. thanks anyway.. btw i knew this method.. i did it by this but the ques states do it by substitution.. @ganeshie8 can u please help me??

OpenStudy (solomonzelman):

maybe sub in u for e^x ?

OpenStudy (solomonzelman):

The you would be getting the same but with a 'u'\[\int\limits_{ }^{ } u^{-2}~du+\int\limits_{ }^{ } u^{2}~du+\int\limits_{ }^{ } 2~du\]

OpenStudy (solomonzelman):

Then*

OpenStudy (aayushi.somani):

yeah there no difference.. if i substitute 1/e^x then?

OpenStudy (solomonzelman):

basically, no

OpenStudy (aayushi.somani):

then?

OpenStudy (solomonzelman):

I mean "then" after you substitute `u` for e^x, you would be getting the same thing, except that you are wasting a step by doing that.

OpenStudy (aayushi.somani):

and what if i substitute 1/e^x??

OpenStudy (solomonzelman):

Well, basically i it will be like this, (using § for an integral sign) § (e^x + 1/e^x)² dx § (u + 1/u)² du § (u + u\(\scriptsize\color{slate}{^{-1}}\) )² du § (u² + u\(\scriptsize\color{slate}{^{-2}}\)+2 ) du § u² du + § u\(\scriptsize\color{slate}{^{-2}}\) du + § 2 du

OpenStudy (solomonzelman):

the rest you can do yourself, and sub e^x for you when you are done... let me know if you need more help

OpenStudy (aayushi.somani):

ohk.. thanks (:

OpenStudy (solomonzelman):

yw